OpenStudy (anonymous):

~ Joseph is twice as old as Allan (I get this its 2x assuming x is Allan's age) in five years allan's age will be two thirds of Joseph's age. how old are they now? ~the sum of the digits of a two digit number is 12. if the digits are reversed, the new number is 12 less than twice the original. find the new number. ~ the length of a rectangle is 4 less than twice its width. if the perimeter is at most 70 cm, find the biggest possible area of the rectangle. ** I have trouble interpreting word problems :(

OpenStudy (amistre64):

words are always a problem; but its good to just take it piece by piece and construct a math sentence to match

OpenStudy (amistre64):

Joseph is twice as old as Allan: j = 2a in five years allan's age will be two thirds of Joseph's age 2(j+5)/3 = 2a + 5 how old are they now? with any luck i read that correctly ;)

OpenStudy (amistre64):

we can say the j=2a or a=j/2; ima use a=j/2 and substitute it into the second part 2(j+5)/3 = 2a + 5 2(j+5)/3 = 2j/2 + 5 expand it all out to see it in a better light 2j/3 + 10/3 = j + 5 collect the js to one side and push everything else to the other side 2j/3 - j = 5 - 10/3 factor out a j j (2/3 - 1) = 5 - 10/3 and divide of the extra j = (5 - 10/3)/(2/3 - 1) = -5 .... what that tells me is that either my setup was incorrect, or j is very special :)

OpenStudy (anonymous):

thank you!!! thank you!! at least I understand how this thing goes now :D I still have a bit of a problem translating word problems into equations though, but thanks!!

OpenStudy (amistre64):

2(2a+5)/3 = 2a + 5 (4a+10)/3 = 2a + 5 a(4/3 - 2) = 5 - 10/3 a = (5 - 10/3)/(4/3 - 2) = -2.5 yeah, i still have issues with them as well .....

OpenStudy (amistre64):

another way to approach this; is to consider that in 5 years, if we want to keep things as nice integers; in 5 years joseph has to be a multiple of 3 3 6 9 12 15; 15 looks kinda 5ish to me 15 - 5 = 10, and 10/2 is 5 j = 2a 10 = 2(5) add 5 years 15 = 5.5.5; 10 = 5.5 which is 2/3 of it